# Expanding Deep Learning-based Sensing Systems with Multi-Source Knowledge Transfer

### v1

@@ Abstract.
Expanding the existing sensing systems to provide high-quality deep learning models for more domains, such as new users or environments, is challenged by the limited labeled data and the data and device heterogeneities. While knowledge distillation methods could overcome label scarcity and device heterogeneity, they assume the teachers are fully reliable and overlook the data heterogeneity, which prevents the direct adoption of existing models. To address this problem, this paper proposes an efficient knowledge transfer framework, HaKT, to expand sensing systems. It first selects multiple high-quality models from the system at a low cost and then fuses their knowledge by assigning sample-wise weights to their predictions. Later, the fused knowledge is selectively injected into the customized models for new domains based on the knowledge quality. Extensive experiments on different tasks, modalities, and settings show that HaKT outperforms stat-of-the-art baselines by at most 16.5% accuracy and saves up to 39% communication traffic.
Sensing System, Deep Learning, Transfer Learning, Knowledge Distillation
copyright: acmcopyrightjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: ; July 2017; Washington, DC, USAprice: 15.00isbn: XXX-X-XXXX-XXXX-X/XX/XX
1. Introduction
The integration of deep learning with sensing systems has gained significant attention in recent years due to its effectiveness in processing diverse sensory data across a wide range of applications (Xu et al., 2023; Ouyang et al., 2022, 2023; Ji et al., 2023). As deep learning techniques continue to mature and computing resources advance, there is an increasing demand for the large-scale deployment of these models in sensing systems (Doe, 2023). However, despite this growing need, the challenge of expanding these deep learning-based systems remains largely unaddressed. In this context, expansion refers to the ability to develop effective models that can seamlessly accommodate new users, devices, environments, or datasets.
We identify the problem of expansion in sensory applications as particularly challenging. First, labeling sensor data is both time-consuming and costly, resulting in limited labeled data for new targets, which undermines the effectiveness of direct supervised training methods (Xu et al., 2021). Second, data collected from various sources, such as different users, devices, or environments, often exhibit varying distributions. In more extreme cases, data from different sources may represent different categories, leading to performance degradation or rendering existing models inapplicable when transferred to new targets (Xu et al., 2023). Third, devices across domains may differ in computational power and memory resources, imposing constraints on model architecture and making current models less suitable for new devices (Wen et al., 2023). Moreover, developing models for new targets must also account for overall training and communication costs. Given these challenges, this work addresses a critical question: How can we effectively and efficiently expand existing sensing systems?
Figure 1. Expanding sensing systems is challenging due to label scarcity and large heterogeneity.
@@ 1. Introduction
Figure 1 illustrates a realistic scenario for the sensing systems expansion. Existing domains, such as different users, devices, datasets, or organizations, maintain heterogeneous models for processing local data, referred to as the source domains. Due to privacy concerns, the data from one source domain cannot be accessed by others. The new targets, i.e., target domains, have limited labeled data but many unlabeled data due to the large labeling overhead. The data within these domains are non-independent and identically distributed (non-IID). The objective is to provide high-quality customized models for the target domains at a minimal cost.
Existing approaches struggle to work in the practical scenario. Directly training models on target domain data is less effective due to label scarcity (Xu et al., 2021; Song et al., 2022). Federated learning frameworks train models using distributed data from multiple clients (Ouyang et al., 2021; Zheng et al., 2023; Cai et al., 2023). However, many of them provide one unified model for all clients and may incur frequent model transmissions and heavy retraining overhead once new targets join. Transfer learning aims to apply existing models to new domains with data or task heterogeneity (Pan and Yang, 2009; Tan et al., 2018; Tong et al., 2021; He et al., 2023). To handle data heterogeneity, domain adaptation techniques are widely used to enhance model robustness by aligning feature distributions across domains (Zhu et al., 2020; Wilson et al., 2021; He et al., 2023). Nevertheless, transfer learning works overlook device heterogeneity, making it difficult to adopt existing models across diverse hardware environments. Knowledge distillation addresses device-side constraints by transferring knowledge from teacher models to student models (Gou et al., 2021; You et al., 2017; Hinton et al., 2015; Borup et al., 2023). Yet existing methods assume that the teacher models are fully reliable and do not account for the impact of data heterogeneity. The non-IID data across domains makes source models less accurate on target domains. Therefore, there is a significant gap in addressing the system expansion problem.
To address this gap, we notice that while source models are trained on non-IID data, they may still retain valuable ”knowledge”, such as the ability to interpret sensor characteristics and distinguish between different classes. Our key idea is to leverage the knowledge from source models to assist target domains, which can help mitigate label scarcity. However, due to data heterogeneity, the knowledge from certain source models may be ineffective in target domains or result in conflicting predictions for the same sample—an issue we refer to as knowledge conflicts in this paper. Therefore, the central challenge is: How can we transfer knowledge from source models to train customized target models? This involves addressing three specific subchallenges.
First, it is difficult to identify models with high-quality knowledge, as their performance in the source domains does not necessarily reflect their effectiveness in target domains. Besides, evaluating these models on target domain data is problematic due to label scarcity. Second, effectively utilizing conflicting knowledge from different sources to train target models is challenging. Resolving knowledge conflicts requires prioritizing certain models, but the reliability of knowledge from each model can vary dynamically across different samples and domains. Third, minimizing system overhead during the expansion process is essential. Evaluating every source model on target domains can incur significant communication and execution costs, especially when there are many large models. Moreover, training target models with multiple source models can be expensive and time-consuming, necessitating optimization to reduce overhead.
To address these challenges, a practical Heterogeneity-aware Knowledge Transfer framework (HaKT) is proposed: 1) To identify high-quality source models, HaKT first utilizes simple statistical features, such as mean values, to filter out source domains with low similarity to the target domain. The models from the remaining domains are then further selected based on their performance on both labeled and unlabeled data in the target domain. 2) To address knowledge conflicts, an attention-based mixer is trained to assign sample-wise weights to each model’s predictions by measuring the similarity between the representations extracted by the selected models and the target model. A knowledge dictionary is constructed to selectively store the fused predictions, considering their fluctuating quality during training. This stored knowledge is later injected into the target model along with the ground-truth labels, with the transfer speed dynamically adjusted according to the quality of the knowledge. 3) To minimize system overhead, the model selection process is encapsulated in a communication protocol that only transmits models with high potential (judged based on statistical features) for further evaluation in the target domains. A low-cost joint training scheme is implemented to simultaneously update the mixer, the target model, and the partially frozen source models. To reduce encoding overhead, the mixer fuses predictions for all the data to update the knowledge dictionary only when its performance improves.
Extensive experiments have been conducted across various modalities and tasks, including Human Activity Recognition, Gesture Recognition, and Image Classification. Compared with five state-of-the-art baselines, HaKT achieves up to a 16.5% improvement in accuracy and reduces communication traffic by up to 39% during sensing systems expansion. The key contributions are summarized as follows:
  1. (1)
    We address a practical sensing system expansion problem characterized by label scarcity and both data and device heterogeneities.
  2. (2)
    We propose a general knowledge transfer framework, HaKT, which efficiently selects, fuses, and injects knowledge to deliver high-quality customized models for targets while maintaining practical system overhead.
  3. (3)
    We comprehensively evaluate the proposed framework across various tasks, modalities, architectures, and settings, demonstrating superior performance compared with state-of-the-art baselines at a reduced cost.
The paper is organized as follows: Section 2 presents the motivation. Section 3 provides the framework overview. Section 4 describes the technical details. Sections 5 show the experiment results. Section 6 presents discussions. Section 7 describes the related works. Section 8 concludes the paper.
@@
Figure 2. Motivation study. (a) When label scarcity is presented, direct training is less effective compared with transferring high-quality source models. The $\gamma$ is set to 10%. (b) Existing methods cannot fully leverage the knowledge from source domain models. (c) The knowledge conflicts between source models are hard to be resolved by existing knowledge fusion methods. One bar refers to the prediction result on one sample.
2. Motivation Study
2.1. Problem Formulation
There are $N_{S}$ source domains $\mathcal{D}_{S}$ within an sensing system, each possessing heterogeneous models for data analytics, represented by $\{\{f_{i},g_{i}\},i=1,\ldots,N_{S}\}$, where $f_{i}$ and $g_{i}$ are the encoder and classifier for the $i$-th domain, respectively. The objective is to expand the sensing system to $N_{T}$ target domains, $\mathcal{D}_{T}$, by providing high-quality customized models for each target. However, only $\gamma\%$ of the data in the target domains are labeled, while the remainder remains unlabeled due to the high cost of labeling. Besides, the raw data in the source domains are inaccessible due to privacy constraints.
2.2. Potential of Knowledge Transfer
To demonstrate the potential of knowledge transfer for providing high-quality models, we compare two naive approaches: direct training and model transfer. Direct training only train with the limited labeled data (10%) in the target domain, whereas the model transfer applies the source models to the target domain. Both methods are evaluated on the HARBox dataset (see Section 5.1). In Figure 2(a), the performance of the source models before transfer (after transfer) is measured on the source (target) domain test sets.
Figure 2(a) demonstrates that direct training yields only around 50% accuracy under conditions of label scarcity. When models from the source domains are applied to the target domain data, they suffer varying degrees of performance degradation due to data heterogeneity. However, the model from domain 5 maintains an accuracy of over 60%, which might be due to a more similar data distribution. These results suggest that transferring high-quality models with appropriate selection strategies to new targets could outperform direct training. However, device heterogeneity and the customized requirements of target domains may restrict the direct adoption of these models. Additionally, when source and target domains exhibit substantial differences in data distribution or contain different categories, even the best-performing model may still fall short of the desired accuracy.
2.3. Limitations of Existing Methods
We further investigate the limitations of existing knowledge transfer methods on the HARBox dataset, including one domain adaptation method (LEAD (Qu et al., 2024)), one model merging method (MEHLSoup (Li et al., 2024)), and one knowledge distillation method (DistillWeighted (Borup et al., 2023)). See Section 5.1 for detailed descriptions. Additionally, we introduce a variant of DistillWeighted, named Random, where source models are randomly selected during the knowledge transfer process. The target models are configured as TPN-M (see Section 5.1), and we report the average accuracy across ten target domains.
In Figure 2(b), both LEAD and MEHLSoup demonstrate suboptimal performance, as they are restricted to leveraging knowledge only from source models that match the target model architecture —a limitation that also applies to other domain adaptation and model merging methods. In contrast, knowledge distillation methods like DistillWeighted are not bound by source model architecture, leading to better performance. Moreover, while domain adaptation typically relies on a single source model, utilizing multiple source models, as demonstrated by DistillWeighted, results in better target models. This improvement arises because the knowledge from multiple teacher models is more comprehensive and can compensate for each other (You et al., 2017). However, existing multi-source distillation methods cannot effectively handle the knowledge conflicts among different source models, which negatively impact the quality of the transferred knowledge.
Knowledge conflicts arise when heterogeneous source models produce different predictions for the same input. Figure 2(c) illustrates the per-sample predictions of two models selected by DistillWeighted, which differ significantly. DistillWeighted simply assign fixed weights to each model, which fails to resolve these conflicts. When some selected models are inaccurate on the target due to data heterogeneity, the accuracy of the fused predictions can be even lower than that of the best single model.
3. Framework Overview
We further analyze the challenges in expanding sensing systems and introduce the proposed framework, HaKT.
3.1. Challenges in Scaling Sensing Systems
@@
To efficiently expand sensing systems to new target domains, three key challenges must be addressed:
@@
First, it is difficult to identify high-quality source models within the sensing system. The performances of a model in its source domain does not necessarily reflect its performance after the transfer due to the varying discrepancies between different source and target domain pairs. Estimating performance degradation based on discrepancy measurements, such as Maximum Mean Discrepancy (Shlens, 2014; Chang et al., 2020), is impractical because these metrics require access to both source and target domain data, which violates data privacy constraints. Additionally, evaluating model performance solely on the limited labeled data in the target domains may not accurately reflect the model’s in-the-wild performance, as those data can hardly capture the diversity of the target data [C1].
@@
Second, transferring knowledge from multiple source models to customized target models is challenging. Due to data heterogeneity, the knowledge in source models may not be directly applicable to the target domain data. Additionally, knowledge conflicts occur when source models make conflicting predictions for the same input, leading to contradictory update directions and less effective model training. Resolving these knowledge conflicts is particularly difficult because the quality and relevance of the knowledge from each model can vary dynamically across different samples. For example, a model’s prediction may be accurate for some subsets of target data but unreliable for others. This context-dependent variability makes it challenging to determine which model’s knowledge should be prioritized [C2].
@@
Third, minimizing system overhead is crucial during expansion. Evaluating source models on the target domains would incur significant communication and execution costs, especially given the large number of source and target domains in sensing systems. However, if all models are not tested on the target domains, some high-quality models might be discarded, leading to suboptimal performance for new targets. Furthermore, leveraging the knowledge from multiple models to train the target model can also introduce large training overhead, which needs to be minimized particularly for resource constraint devices [C3].
Figure 3. Overview of the HaKT framework.
3.2. System Architecture
@@
To facilitate efficient sensing system expansion, HaKT is proposed to address the limitations of existing methods and overcome the three aforementioned challenges. The overview of HaKT is presented in Figure 3. HaKT first determines a suitable model skeleton either directly provided by the target or from a model library with device profiling. The model library includes open-source models from the Internet and the architectures from the source domains. Then the Efficient Model Selection Protocol is evoked to select high-quality source models at a low cost. Later, the Sample-wise Knowledge Fusion is performed to aggregate the conflicting knowledge. Subsequently, the target model is trained with the Adaptive Knowledge Injection based on a low-cost training scheme. In HaKT, three core technical components are designed:
  1. (1)
    Efficient Model Selection Protocol involves a two-stage process to identify high-quality models at a low cost. To avoid full model transmission and execution, a coarse-grained selection first filters out less effective models based on lightweight features [C1, C3]. The remaining model candidates are then transmitted to the targets for a fine-grained selection, which evaluates them using both labeled and unlabeled data [C1].
  2. (2)
    Sample-wise Knowledge Fusion is proposed to resolve the knowledge conflicts. An attention-based mixer is trained to assign sample-specific weights to the predictions of the selected models based on the representation proximity extracted by the selected source and target models [C2]. Additionally, the selected models are partially frozen and fine-tuned to enhance the quality of the fused knowledge [C2, C3].
  3. (3)
    Adaptive Knowledge Injection facilitates the injection of fused knowledge to the target models. It utilizes a Knowledge Dictionary to selectively store the fused knowledge and an adaptive learner to dynamically adjust the importance of the knowledge based on its quality during training [C2]. To minimize training overhead, the system only fuses predictions for all the data to update the Knowledge Dictionary after the performance of the mixer shows improvement [C3].
HaKT is a highly flexible and decentralized framework for sensing system expansion. Without constraints in training methods, architectures, or data quality, any domains could be included as source domains to contribute knowledge for new targets. Moreover, the constructed target models can be further included as source models for other new targets. Besides, the model customization process for the new targets in HaKT is cheap in terms of communication and training overhead. Section 4 further introduces the details of HaKT on providing a customized model for one target, the processes of which are scalable and repeatable for any number of targets.
4. Design of HaKT
4.1. Efficient Model Selection Protocol
The process of the Efficient Model Selection Protocol is detailed in Figure 4. In the Feature-based Coarse Selection stage (Steps FbCS.1-3), domains with potentially high-performing models are identified by comparing lightweight features between the source and target domains. Only source domains with high feature similarity transmit their models to the target, reducing communication traffic and model execution costs. During the Centroids-Accuracy Joint Selection stage (Steps CAJS.1-3), the transmitted models extract centroids and estimate their accuracy on the target domain data, which are leveraged to select the final models.
4.1.1. Feature-based Coarse Selection.

### v2

# Expand Heterogeneous Learning Systems with Selective Multi-Source Knowledge Fusion

@@ Abstract
Expanding existing learning systems to provide high-quality customized models for more domains, such as new users, is challenged by the limited labeled data and the data and device heterogeneities. While knowledge distillation methods could overcome label scarcity and device heterogeneity, they assume the teachers are fully reliable and overlook the data heterogeneity, which prevents the direct adoption of existing models. To address this problem, this paper proposes a framework, HaT, to expand learning systems. It first selects multiple high-quality models from the system at a low cost and then fuses their knowledge by assigning sample-wise weights to their predictions. Later, the fused knowledge is selectively injected into the customized models based on the knowledge quality. Extensive experiments on different tasks, modalities, and settings show that HaT outperforms state-of-the-art baselines by up to 16.5% accuracy and saves up to 39% communication traffic.
Deep Learning Systems, Transfer Learning, Knowledge Distillation
1 Introduction
Deployment of customized learning models presents a critical dilemma. While building specialized models for each user, device, or environment (domain) on the edge can yield fine-grained performance and preserve privacy (Lu et al., 2024; Kong et al., 2023), producing these per-domain models is expensive. Each customized model demands substantial labeling and training efforts, yet in practice, many domains have only scarce labeled data due to prohibitive labeling costs (Gao et al., 2024; Dai et al., 2024). Moreover, modern learning systems typically serve many such domains (potentially in the hundreds or thousands), in which cases the labeling and retraining overhead becomes unsustainable.
While one might consider transferring existing models to new domains (Phan et al., 2024; Lu & Sun, 2024), significant challenges arise from data and device heterogeneity. Models trained on a particular data distribution may fail to generalize to another, suffering from performance drops or inapplicability to unseen categories. Additionally, resource constraints of different devices, such as memory or computational power, further complicate direct model reuse (Li et al., 2024b). Consequently, a more systematic strategy is required to expand existing pool of models to accommodate new target domains without incurring massive cost. In this work, we refer to this process as learning system expansion.
@@ 1 Introduction
Figure 1 illustrates a realistic scenario for the learning systems expansion. Existing domains, such as different users, devices, datasets, or organizations, maintain heterogeneous models for processing local data, referred to as the source domains. The new targets, i.e., target domains, have limited labeled data but many unlabeled data due to the large labeling overhead. The data within these domains are non-independent and identically distributed (non-IID) with potential shifts in label space. We pose a critical question: How can we effectively and efficiently expand learning systems?
Figure 1: Expanding learning systems is challenging due to label scarcity and large heterogeneity.
@@
Figure 2: Motivation study. (a) When only 10% data is labeled, direct training is less effective compared with transferring high-quality source models. (b) While leveraging more source models may lead to better performance, existing methods cannot fully exploit the knowledge from source domain models. (c) Knowledge conflicts between source models are hard to be resolved by existing knowledge fusion methods. One bar refers to the prediction result on one sample.
Existing approaches struggle to handle this question. To handle data heterogeneity, domain adaptation is widely studied to enhance model robustness by aligning feature distributions across domains (Wilson et al., 2021; He et al., 2023; Qu et al., 2024). Nevertheless, those works overlook device heterogeneity, making it difficult to adopt existing models across diverse hardware environments. Knowledge distillation addresses device-side constraints by transferring knowledge from teacher models to student models (Hinton et al., 2015; Gou et al., 2021; Borup et al., 2023; Peng et al., 2024). Yet existing methods assume that the teacher models are fully reliable and do not account for the impact of data heterogeneity. The non-IID data across domains makes source models less accurate on target domains. Therefore, there is a significant gap in addressing the expansion problem.
To address this gap, we conduct a motivation study and observe that, while each source model is trained on potentially non-IID data, it can still retain valuable “knowledge”—as illustrated by model 5 in Figure 2(a)—such as the ability to interpret data characteristics and distinguish between classes. Moreover, as shown in Figure 2(b), combining multiple source models can yield more comprehensive knowledge that compensates the weaknesses of individual models. Based on those observations, our key idea is to exploit the high-quality knowledge from multiple source models to train customized target models.
However, not all knowledge is beneficial for target domains. As shown in Figure 2(b), blindly combining all available models may result in suboptimal performance. Yet identifying high-quality source models is challenging because performance in source domains does not necessarily indicate effectiveness in the target domain, and label scarcity limits direct evaluation. Furthermore, knowledge from different sources may conflict, as illustrated in Figure 2(c). Resolving these conflicts requires prioritizing specific models, yet the reliability of each model’s knowledge can vary dynamically across samples and domains. Additionally, selecting high-quality source models to train target models can be expensive, making optimization essential to reduce overhead. Overall, selectively aggregating and injecting multi-source knowledge into target models while maintaining reasonable overhead is crucial for learning system expansion.
To address these challenges, a practical Heterogeneity-aware Knowledge Transfer framework (HaT) is proposed: 1) To identify high-quality source models, HaT first utilizes simple statistical features to filter out low-quality source domains. The remaining models from are then further selected based on their performance on the target domain data. 2) To address knowledge conflicts, an attention-based mixer is trained to assign sample-wise weights to each model’s predictions by measuring the representations similarity between models. A knowledge dictionary is constructed to selectively store the fused predictions. This stored knowledge is later injected into the target model, with the transfer speed dynamically adjusted based on the quality of the knowledge. 3) To minimize system overhead, the model selection process is encapsulated in a communication protocol that only transmits models with high potential for further evaluation in the target domains. A low-cost joint training scheme is implemented to simultaneously update the mixer, the target model, and the partially frozen source models.
Extensive experiments have been conducted across various modalities and tasks, including Human Activity Recognition, Gesture Recognition, and Image Classification. The results show that HaT achieves up to 16.5% higher accuracy, which also reduces communication traffic by up to 39% and source model execution overhead by 40%. The key contributions are as follows:
1) We address a practical learning system expansion problem characterized by label scarcity and both data and device heterogeneities.
2) We propose a general system expansion framework, HaT, which efficiently selects, fuses, and injects multi-source knowledge to deliver high-quality customized models for targets with practical system overhead.
3) We evaluate the framework across various tasks, modalities, and architectures, demonstrating superior performance compared with baselines.
2 Related Works
Transfer Learning. Transfer learning explores methods to apply source models for new targets, addressing data or task heterogeneities (Pan & Yang, 2009; Tan et al., 2018). In particular, domain adaptation has been extensively studied to align feature distributions between domains (Zhu et al., 2020; Wilson et al., 2021; He et al., 2023; Qu et al., 2024). However, these approaches typically require access to both source and target domain data, which may not be feasible. In contrast, test-time adaptation techniques adapt models using only test data, enabling continual learning (Gong et al., 2024; Karmanov et al., 2024). Additionally, multi-source transfer learning methods aim to select source models with better generalizability (Tong et al., 2021; Agostinelli et al., 2022). Despite these advancements, most approaches fail to address device heterogeneity, which might hinder the direct application of source models to target domains.
Knowledge Distillation. In knowledge distillation, an efficient student model is trained using the knowledge from one or more teacher models, such as their predicted pseudo labels or intermediate features (Hinton et al., 2015; Liu et al., 2020; Vemulapalli et al., 2024; Peng et al., 2024). Specifically, multi-teacher distillation approaches (Borup et al., 2023; Liu et al., 2020; Zhang et al., 2022) aggregate the knowledge of multiple teachers by assigning weights, aiming to provide the student model with more accurate and comprehensive knowledge. Most knowledge distillation studies focus on a closed-set problem, where high-quality teacher models are predetermined and readily available (Hinton et al., 2015; Liu et al., 2020; Zhang et al., 2022; Borup et al., 2023). However, in the learning system expansion problem, the knowledge from source domain models may not directly transfer to the target domain due to data heterogeneity, leading to suboptimal performance.
Model Customization. Model customization has been extensively studied to meet specific computational and performance requirements (Wen et al., 2023; Li et al., 2024b). Some works explore pre-deployment or post-deployment model generation techniques (Cai et al., 2020; Wen et al., 2023) to search optimal architecture in terms of latency and accuracy. In contrast, HaT emphasizes the knowledge transfer process from the selected source models to any target models that satisfy the customized needs of target domains.
3 Heterogeneous Learning Systems Expansion
3.1 Problem Formulation
We define the Multi-Round System Expansion (MRSE) problem to address the ever-growing nature of learning systems. At round $j$, there are $N^{S}(j)$ existing source domains, denoted as $\mathcal{D}^{S}(j)=\{\mathcal{D}^{S}_{i}(j)\}_{i=1}^{N^{S}(j)}$, and $N^{T}(j)$ target domains, $\mathcal{D}^{T}(j)=\{\mathcal{D}^{T}_{i}(j)\}_{i=1}^{N^{T}(j)}$. Each target domain in $\mathcal{D}^{T}(j)$ requires high-quality, customized models to meet its unique requirements. Once the models for target domains $\mathcal{D}^{T}(j)$ are curated, these domains become source domains in the subsequent round: $\mathcal{D}^{S}(j+1)=\mathcal{D}^{S}(j)\cup\mathcal{D}^{T}(j)$. The knowledge in $\mathcal{D}^{S}(j+1)$ is then leveraged to curate models for target domains in $\mathcal{D}^{T}(j+1)$. The primary objectives are: 1) maximize performance of the curated models on target domainsp; 2) minimize curation overhead, which includes communication and computation costs.
To better understand the process, the MRSE problem can be decomposed into individual One-Time System Expansion (OTSE) problems. For a specific target domain $\mathcal{D}^{T}_{i}(j)=\{X^{T}_{i},Y^{T}_{i},\zeta^{T}_{i}\}$, the goal is to curate a model based on the knowledge from source domains $\mathcal{D}^{S}(j)=\{X^{S}_{i},Y^{S}_{i},NN^{S}_{i},\zeta^{S}_{i}\}$. However, source and target domains exhibit distributional differences between $X^{S}_{i}$ and $X^{T}_{i}$, which hinder the direct applicability of the source model $NN^{S}_{i}=\{f^{S}_{i},g^{S}_{i}\}$. The $f^{S}_{i}$ and $g^{S}_{i}$ are the encoder and the classifier. The label sets of source domains $Y^{S}_{i}$ and target domain $Y^{T}_{i}$ may not fully overlap, introducing additional complexity during expansion. Moreover, target domains impose constraints $\zeta^{T}_{i}$, including memory usage and inference speed requirements, which must be considered during model curation. Besides, due to the high cost of labeling, only few ($\gamma\%$) data in $\mathcal{D}^{T}_{i}(j)$ is labeled, which is a general assumption to handle potential label space difference.
3.2 Technical Challenges
@@
To expand learning systems to new target domains, three key challenges must be addressed:
@@
First, it is difficult to identify high-quality source models within the systems. Estimating performance degradation caused by the data discrepancies based on discrepancy measurements is less practical if requiring access to both source and target domain data (Chang et al., 2020; Qu et al., 2024). Additionally, evaluating models solely with few labeled data in the target domains may not accurately reflect the model’s in-the-wild performance, as those data can hardly capture the diversity of the target data.
@@
Second, transferring knowledge from multiple source models to customized target models is challenging due to the knowledge conflicts. Source models make conflicting predictions for the same input, which leads to contradictory update directions and less effective model training. Resolving knowledge conflicts is difficult because the quality and relevance of the knowledge from each model can vary dynamically across different samples. This context-dependent variability makes it challenging to determine which model’s knowledge should be prioritized.
@@
Third, minimizing system overhead is crucial during expansion. Evaluating source models on the target domains would incur significant communication and execution costs, given the large number of domains in the systems. However, if all models are not tested in the target domains, high-quality models might be discarded, leading to suboptimal performance. Furthermore, leveraging the knowledge from multiple models to train the target model can also introduce large training overhead, which needs to be minimized.
4 Design of HaT
4.1 Framework Overview
@@
HaT is designed to address the three aforementioned challenges, the process of which is presented in Figure 3. HaT first invokes the Efficient Model Selection Protocol to select high-quality source models at a low cost. Later, the Sample-wise Knowledge Fusion is performed to aggregate the conflicting knowledge. Subsequently, the target model is trained with the Adaptive Knowledge Injection based on a low-cost training scheme. We further introduce the details of providing a customized model for one target $\mathcal{D}_{t}$ (instead of $\mathcal{D}_{t}^{T}$ for clarity), the processes of which are scalable and repeatable for any number of targets.
Figure 3: Framework overview of HaT.
4.2 Efficient Model Selection Protocol
The Efficient Model Selection Protocol aims to select high-quality source domain models at a low communication and model inference cost with two stages of selection: